A uniform magnetic field is at right angles to the plane of a wire loop. The field decreases by 0.16 Tin 3.0×10−3 s and the magnitude of the average emf induced in the loop is 79 V .

A) What is the area of the loop? (in m^2)

B) What would be the value of the average induced emf if the field change was the same but took twice as long to decrease? (in Volts)
Answer #1

delta phi = A * delta B

delta phi = A * 0.16

e m f = delta phi / delta t

79 = (A * 0.16) / 0.003

A = 1.48 m2

---------------------------------------------------

e m f = delta phi / delta

= (A * 0.16) / 0.006

= (1.48 * 0.16) / 0.006

= 39.47 V